# Graph Representation of Narrative Context: Coherence Dependency via Retrospective Questions

### v1

@@ A.2.1 Case1
Current Context:
However, Oberyn stands too close to his seemingly defeated opponent, and Gregor manages to trip and seize him. Berserk with fury, Gregor grabs Oberyn by the throat and lifts him off the ground, smashing out most of his teeth with a single devastating punch. Climbing on top of Oberyn, Gregor finally admits for all to hear that he raped and killed Elia as he gouges out Oberyn’s eyes with his thumbs before crushing the Viper’s skull between his hands, which he proclaims having done the same to his sister. As Ellaria screams in horror, a stunned silence sweeps over the crowd. The short joyful moments for Tyrion and Jaime are shattered, as Tywin stands and proclaims the will of the gods is clear: Tyrion is guilty and sentenced to death. Tyrion cannot even reply, shockingly staring in catatonic astonishment at Oberyn’s skull-crushed corpse, as does Jaime; the only different reaction is from Cersei, who stares at Oberyn’s slaughtered body, listening to Tyrion’s death sentence while smirking in vindication.
Prior Context:
@@
Generated Question (Valid):
@@
Generated Question (Invalid):
@@ A.2.2 Case2
Generated Question (Valid):
@@ A.2.3 Case3
Generated Question (Valid):
@@
Generated Question (Invalid):
@@ A.2.4 Case4
Generated Question (Invalid):
@@ A.2.5 Case5
Geneated Question (Invalid):

### v2

@@ A.2.1 Case1
Current Context: {adjustwidth}-0.4cm-0.4cm
At Craster’s Keep, Locke scouts the keep for the party of the Night’s Watch sent to eliminate the traitors holed up there; in his reconnaissance, Locke finds the hut where Bran Stark, Jojen, Meera and Hodor are being held captive. Reporting back to Jon Snow and the others, Locke tells them that only eleven traitors are present and most of them are drunk and won’t prove much of a threat. He also lies about the hut where Jon’s brother Bran and his group are being imprisoned, claiming there are only hounds kept inside and that they should keep away from it to prevent the dogs alerting their enemy. Believing Locke, Jon agrees and tells the party they attack at nightfall.
Prior Context: {adjustwidth}-0.4cm-0.4cm
@@
Generated Question (Valid): {adjustwidth}-0.4cm-0.4cm
@@
Generated Question (Invalid): {adjustwidth}-0.4cm-0.4cm
@@ A.2.2 Case2
Generated Question (Valid): {adjustwidth}-0.4cm-0.4cm
@@ A.2.3 Case3
Generated Question (Valid): {adjustwidth}-0.4cm-0.4cm
@@
Generated Question (Invalid): {adjustwidth}-0.4cm-0.4cm
@@ A.2.4 Case4
Generated Question (Invalid): {adjustwidth}-0.4cm-0.4cm
@@ A.2.5 Case5
Geneated Question (Invalid): {adjustwidth}-0.4cm-0.4cm